# Adaptive Visual Scene Understanding: Incremental Scene Graph Generation

### v1

@@ 5 Discussion
In the dynamic world, the incremental introduction of new objects and new relationships in scenes presents significant challenges for scene graph generation (SGG) models to effectively adapt and generalize without forgetting previously acquired semantic knowledge. However, despite the progress made in SGG and continual learning research, there remains a lack of comprehensive investigations specifically targeting the unique challenges of Continual Scene Graph Generation (CSEGG). To close this research gap, we take the initial steps of operationalizing CSEGG and introducing benchmarks, datasets, and evaluation protocols. Our study delves into three distinct learning scenarios, thoroughly examining the interplay between continual object detection and continual relationship classification for existing CSEGG methods under long-tailed class-incremental settings.
Our experimental results reveal intriguing insights. First, applying standard continual learning approaches combined with long-tailed techniques to SGG models yields moderate improvements. However, a notable performance gap persists between current CSEGG methods and the joint training upper bound. Second, we investigated the model’s generalization ability and found that the models are capable of generalizing to classify known relationships involving unfamiliar objects. Third, we compared the CSEGG performance of the traditional CNN-based and the transformer-based SGG models as backbones. We observed consistent relative CSEGG performance across all continual learning methods using both backbones, with CNN-SGG models underperforming SGTR-based ones.
Moving forward, there are several key avenues for future research. Our current endeavors focus on learning CSEGG problems from static images in Independent and Identically Distributed (i.d.d) manner, diverging from human learning from video streams. Future research can look into CSEGG problems on video streams. Our plans also involve expanding continual learning baselines and integrating more long-tailed distribution sampling techniques. Furthermore, we aim to construct a synthetic SGG dataset to systematically quantify the aspects of SGG that influence continual learning performance under controlled conditions. Although the CSEGG method holds promise for many downstream applications like monitoring systems, medical imaging, and autonomous navigation, we should also be aware of its misuse in privacy, data biases, fairness, security concerns, and misinterpretation. We invite the research community to join us in maintaining and updating the safe use of CSEGG benchmarks, thereby fostering its advancements in this field.
Ethics Statement
The development and deployment of Scene Graph Generation (SGG) technology present potential negative societal impacts that warrant careful consideration (Li et al., 2022b). Firstly, privacy concerns arise as SGG may inadvertently capture sensitive information from images, potentially violating privacy rights and raising surveillance issues. Secondly, bias and fairness challenges persist, as SGG algorithms can perpetuate biases present in training data, leading to discriminatory outcomes that reinforce societal inequalities. Misinterpretation and misclassification by SGG algorithms could result in misinformation and incorrect actions, impacting decision-making. The risk of manipulation and misuse of SGG-generated scene representations for malicious purposes is also a concern. For example, attackers might manipulate scene graphs to deceive systems or disrupt applications that rely on scene understanding.
Reproducibility Statement
We are committed to ensuring the reproducibility and transparency of our research. In accordance with the guidelines set forth by ICLR 2024, we provide detailed information to facilitate the replication of our experiments and results.
  1. 1.
    Code Availability: All code used for our experiments is available at here.
  2. 2.
    Data Availability: Any publicly accessible datasets used in our research are specified in the paper, along with their sources and access information.
  3. 3.
    Experimental Details: We have documented the specific details of our experiments, including hyper-parameters, model architectures, and pre-processing steps, to enable others to replicate our results.
We are dedicated to supporting the scientific community in replicating and building upon our work. We welcome feedback and collaboration to ensure the robustness and reliability of our research findings.
Acknowledgement
@@
Appendix A Appendix
A.1 Motivation for Learning Scenarios
@@
Within this section, we present practical applications that demonstrate the real-world relevance of investigating CSEGG across all three learning scenarios.
A.1.1 Scenario 1 (S1): Relationship Incremental Learning
@@
Within medical imaging, an agent must acquire the ability to detect cancerous cells within primary tumors, like colon adenocarcinoma. Subsequently, it must extend this proficiency to identifying the same cell types within metastatic growths that manifest in different bodily regions, such as lymph nodes or the liver. In this instance, the identical cancer cell disseminates to fresh organs or tissues, progressively establishing new relationships with other cells over the course of time.
A.1.2 Scenario 2 (S2): Scene Incremental Learning
@@
The CSEGG model’s capacity to incorporate new objects and new relationships while retaining existing knowledge finds pivotal application in security contexts. Consider a company developing video-based security systems for indoor environments, capturing prevalent indoor objects and relationships. Expanding to outdoor settings like parking lots or restricted compounds demands retraining the model with new outdoor data alongside previous indoor data, ensuring operational effectiveness in both realms. The outdoor context introduces new objects like "cars" and relationships like "driving", distinct from indoor scenarios featuring "chair" and "sitting." Employing CSEGG allows the company to focus on new objects and relationships while retaining indoor insights.
@@
A.1.3 Scenario 3 (S3): Scene Graph Generalization In Object Incremental Learning
A prime example is the ongoing research on deep sea exploration for autonomous navigation systems, where undiscovered flora and fauna reside beneath the ocean’s surface. Encountering new and unidentified species becomes manageable through SGG’s ability to understand spatial relations. The robot discerns the object’s proximity or orientation even without precise identification of the spieces, enhancing its autonomous navigation ability. Likewise, in deep space exploration, SGG aids in recognizing spatial relationships with previously unseen space debris, aiding in path-planning. In essence, SGG’s relationship generalization empowers robots to navigate and plan routes in unfamiliar terrains, such as deep sea and deep space, where novel encounters demand adaptable responses.
A.1.4 Profound Impacts Beyond the Problem of CSEGG
We wish to emphasize that the challenge of continuous Scene Graph Generation (SGG) applied to real-world images exhibits two distinctive attributes. Techniques developed to address these attributes possess the potential to be widely applicable across various domains and subsequent tasks. Firstly, SGG in a continuous learning context frequently deals with data stemming from distributions that exhibit long-tailed characteristics. The methodologies formulated within the realm of Continual SGG can be adapted to address broader long-tailed continuous learning issues, such as classifying bird sounds in ecological studies. Secondly, the process of continual SGG demands the ability to engage in reasoning and integrate knowledge over time. As an illustration, in the context of retail inventory management, merely learning to identify durians (a tropical fruit uncommon in US supermarkets) is insufficient. The model must also cultivate the capability to integrate this new information into its existing knowledge database, ensuring that this new product is positioned alongside other fruits within the grocery store.
A.1.5 Design Motivations for Overlapping Images and Non-overlapping Labels across Tasks
@@
A.2 Implementation Details
A.2.1 SGTR
The SGTR is trained in two stages in a supervised manner. In stage 1, only object detection losses in DETR is (Carion et al., 2020) applied on $O_{i}$. In stage 2, only predicate entity prediction loss is applied on $R_{i}$, which can be further decomposed into L1 and GIOU losses for object/subject/predicate localization (Rezatofighi et al., 2019) and cross-entropy loss for object/subject/predicate classification. In learning scenario S1, we skip Stage 1, and directly load pre-trained weights of DETR for object detection on the entire training set of Visual Genome (Li et al., 2022b). In stage 2 of S1, we freeze the feature extractor, and fine-tune the rest parts of SGTR for predicate entity predictions. As only relationship classes are incrementally introduced in S1, we freeze the entire weights of DETR for detecting all the objects in the scene over tasks. However, empirical results suggest that fine-tuning transformer-based encoders in DETR helps downstream predicate predictions (Li et al., 2022b). Even with fine-tuning DETR in S1, we verify that there is minimal forgetting of detecting all objects in the scene over tasks (see Sec. S14 and Sec. A.4.5). Thus, the forgetting observed in S1 could only be attributed to incremental relationship learning. In Stage 1 of S2 and S3 where object classes are also incrementally introduced over tasks, we load weights of the feature extractor, pre-trained on ImageNet (Deng et al., 2009), and fine-tune the entire DETR (Carion et al., 2020) over all the tasks. Stage 2 of S2 and S3 is the same as S1.
@@ A.2.1 SGTR
A.2.2 CNN-SGG Backbone
Given a scene image $I_{i}$, CNN-SGG utilizes Faster-RCNN(Girshick, 2015) to generate a set of object proposals. The model subsequently extracts visual features of nodes and edges from the set of object proposals. Finally, both edge and node GRUs output a structured scene graph via message passing. The CNN-SGG is trained in two stages in a supervised manner. In stage 1, only object detection losses in Faster-RCNN(Ren et al., 2015) are applied on $O_{i}$. We use the cross entropy loss for the object class and $L1$ loss for the bounding box offsets. In stage 2, the visual feature extractor (VGG-16(Simonyan & Zisserman, 2014) pre-trained on ImageNet (Deng et al., 2009)) and GRUs layers are trained to predict the final object classes, bounding boxes, and relationship predicates using cross-entropy loss and $L1$ loss. In Learning Scenario 1 (S1), similar to the implementation details of SGTR in Sec.3.2, we skip Stage 1, and directly load pre-trained weights of Faster-RCNN for object detection on the entire training set of Visual Genome (Li et al., 2022b). In stage 2 of S1, we load the pre-trained weights of the visual feature extractor (pre-trained on ImageNet) and fine-tune the rest parts of the model. In stage 2 of S1, we load the pre-trained weights of visual feature extractor (pre-trained on ImageNet) and fine-tune the rest parts of the model. In Stage 1 of S2 and S3 where object classes are also incrementally introduced over tasks, we load weights of the Faster-RCNN, pre-trained on ImageNet Deng et al. (2009), and fine-tune it over all the tasks. Stage 2 of S2 and S3 follows the same training regimes as Stage 2 of S1.
@@ A.2.2 CNN-SGG Backbone
A.2.3 Definitions of BWT and FWT Evaluation Metrics

### v3

@@ 6 Discussion
In the dynamic world, adapting scene graph generation (SGG) models to new objects and relationships poses challenges. Despite progress in SGG and continual learning, there is still a gap in understanding Continual Scene Graph Generation (CSEGG). We address this by operationalizing CSEGG, and introducing benchmarks, datasets, and evaluation protocols. Our study explores three learning scenarios, analyzing continual object detection and relationship classification in long-tailed class-incremental settings for CSEGG baselines. Our findings show that integrating sampling methods with CSEGG baselines to address long-tailed distributions moderately eliminates forgetting; however, a large performance gap between current CSEGG baselines and the joint training upper bound persists. To address CSEGG challenges, we propose RAS, a Replays via Analysis by Synthesis method. RAS parses previous task scene graphs into triplet labels for diverse in-context scene graph reconstruction. Based on these re-compositional context-congruent scene graphs, RAS synthesizes images with Stable Diffusion models for replays. Unlike other image replay methods, RAS stores only triplet labels and the model snapshot, maintaining constant memory usage and preserving privacy. Extensive experiments demonstrate our RAS’s superior performance over current CSEGG baselines in knowledge transfers and reducing forgetting. Interestingly, our RAS model is also capable of generalizing to classify known relationships among unfamiliar objects. We discuss ethical concerns in the Sec.0.A.6.
7 Acknowledgement
@@
Appendix 0.A Appendix
0.A.1 Introduction to Three Learning Scenarios
@@
Within this section, we present more details of three learning scenarios and their practical applications.
Figure S1: Label distribution in each task in each learning scenario is presented. In scenario S1 (a) and scenario S3 (c), we use different colors to denote different tasks. The color gradient indicates the frequency of data within a task, with the lighter color denoting the smaller frequency of data in that category. Only the most frequent labels (relationship labels in (a) and object labels in (c)) are provided. See the legend for the total data size per task. In (b) scenario S2 on both objects and relationships, data distributions are presented in the form of small-world networks, where nodes denote object categories and the edges linking object pairs indicate relationships. Thickness in edges implies the diversity of relationships between object pairs. Same color conventions as (a) and (c) are applied. See the legend for triplet sizes.
0.A.1.1 Scenario 1 (S1): Relationship Incremental Learning
While existing continual object detection literature focuses on incrementally learning object attributes [37, 61, 8, 59, 51, 39], incremental relationship classifications are equally important as it provides a deeper and more holistic understanding of the interactions and connections between objects within a scene. To uncover contextual information and go beyond studies of object attributes, we introduce this scenario where new relationship predicates $p_{k}$ are incrementally added in each task (Fig. 2 S1). There are 5 tasks in S1. To simulate the naturalistic settings where the frequency of relationship distribution is often long-tailed, we randomly and uniformly sample relationship classes from head, body and tail categories in Visual Genome [23], and form a set of 10 relationship classes for each task. Thus, the relationships within a task are long-tailed; and the number of relationships from the head categories of each task is of the same scale. To tackle this issue, we allow CSEGG models to see the same images over tasks, but the relationship labels are only provided in their given task (see Sec. 0.A.1.4 for the design motivation). The same reasoning applies in S2 and S3. Example relationship classes from each task and their distributions are provided in Fig. S1.
@@ 0.A.1.1 Scenario 1 (S1): Relationship Incremental Learning
Here, we provide a concrete example application of Scenario 1 in medical imaging. Within medical imaging, an agent must acquire the ability to detect cancerous cells within primary tumors, like colon adenocarcinoma. Subsequently, it must extend this proficiency to identifying the same cell types within metastatic growths that manifest in different bodily regions, such as lymph nodes or the liver. In this instance, the identical cancer cell disseminates to fresh organs or tissues, progressively establishing new relationships with other cells over the course of time.
0.A.1.2 Scenario 2 (S2): Scene Incremental Learning
To simulate the real-world scenario when there are demands for detecting new objects and new relationships over time in old and new scenes, we introduce this learning scenario where new objects $O_{i}$ and new relationship predicates $p_{k}$ are incrementally introduced over tasks (Fig. 2 S2). To select the object and relationship classes from the original Visual Genome [23] for S2, we have two design motivations in mind. First, in real-world applications, such as robotic navigation, robots might have already learned common relationships and objects in one environment. Incremental learning only happens on less frequent relationships and objects. (2) Transformer-based AI models typically require large amounts of training data to yield good performances. Training only on a small amount of data from tail classes often leads to close-to-chance performances. Thus, we take the common objects and relationships from the head classes in Visual Genome as one task, while the remaining less frequent objects and relationships from tail classes as the other task. This results in 2 tasks in total with the first task containing 100 object classes and 40 relationship classes. In the subsequent task, the CSEGG models are trained to continuously learn to detect 25 more object classes and 5 more relationship classes. Same as S1, both the object class and relationship class distributions are still long-tailed within a task (Fig. S1).
Next, we provide two real-world example applications in robot collaborations on construction sites and video surveillance systems.
@@ 0.A.1.2 Scenario 2 (S2): Scene Incremental Learning
The CSEGG model’s capacity to incorporate new objects and new relationships while retaining existing knowledge finds pivotal application in video surveillance contexts. Consider a company developing video-based security systems for indoor environments, capturing prevalent indoor objects and relationships. Expanding to outdoor settings like parking lots or restricted compounds demands retraining the model with new outdoor data alongside previous indoor data, ensuring operational effectiveness in both realms. The outdoor context introduces new objects like "cars" and relationships like "driving", distinct from indoor scenarios featuring "chair" and "sitting." Employing CSEGG allows the company to focus on new objects and relationships while retaining indoor insights.
@@
0.A.1.3 Scenario 3 (S3): Scene Graph Generalization In Object Incremental Learning
We, as humans, have no problem at all recognizing the relationships of unknown objects with other nearby objects, even though we do not know the class labels of the unknown objects. This scenario is designed to investigate whether the CSEGG models can generalize as well as humans. Specifically, there are 4 tasks in total with each task containing 30 object classes and 35 relationship classes. In each subsequent task, the CSEGG models are trained to continuously learn to detect 30 more object classes and learn to classify the same set of 35 relationships among these objects. The class selection criteria for each task follow the same as S1, where the selections occur uniformly over head, body, and tail classes. Example object classes and their label distributions for each task are provided in Fig. S1. Different from S1 and S2, a standalone generalization test set is curated, where the objects are unknown and their classes do not overlap with any object classes in the training set but the relationships among these unknown objects are common to the training set of every task. The CSEGG models trained after every task are tested on the same generalization test sets.
Here, we provide two real-world applications of Scenario 3 in the deep sea and space explorations for autonomous navigation systems.
A prime example is the ongoing research on deep sea exploration for autonomous navigation systems, where undiscovered flora and fauna reside beneath the ocean’s surface. Encountering new and unidentified species becomes manageable through SGG’s ability to understand spatial relations. The robot discerns the object’s proximity or orientation even without precise identification of the species, enhancing its autonomous navigation ability. Likewise, in deep space exploration, SGG aids in recognizing spatial relationships with previously unseen space debris, aiding in path-planning. In essence, SGG’s relationship generalization empowers robots to navigate and plan routes in unfamiliar terrains, such as deep sea and deep space, where novel encounters demand adaptable responses.
0.A.1.4 Design Motivations for Overlapping Images and Non-overlapping Labels across Tasks
@@
0.A.2 Implementation and Training Details
Figure S2: Introduction to backbone SGG models and continual learning baselines. We use Scene graph Generation TRansformer (SGTR) [28] as the backbone SGG model (Sec. 3.2). SGTR consists of four modules indicated by each blue box. Arrows indicate the signal flows among modules. (b) Four continual learning baselines are listed: EWC [22], Replay [48], Naive (Sec. 3.2) and PackNet [38](Sec. 3.2). $\theta^{*}_{A}$ denotes the optimal network parameters after learning on task A. The arrows in colors indicate the shifts of network weights in the parameter space when learning Task B for different baselines.
0.A.2.1 SGTR Backbone
For SGTR as depicted in (Fig. S2 (a)), the approach uniquely formulates the task as a bipartite graph construction problem. Starting with a scene image ($I_{i}$), SGTR utilizes a 2D-CNN and transformer-based encoder to extract image features. These features are then incorporated into a transformer-based decoder, predicting object and subject nodes ($O_{i}$). Predicate nodes ($R_{i}$) are formed based on both image features and object node features, and a bipartite graph ($G_{i}$) is constructed to represent the scene collectively. The correspondence between object nodes ($o_{i}$) and predicate nodes ($r_{k}$) is established using the Hungarian matching algorithm [24]. Experimental results are based on the average over three runs, and the implementation leverages public source codes from [28] and [59] with default hyperparameters.
The SGTR is trained in two stages in a supervised manner. In stage 1, only object detection losses in DETR is [6] applied on $O_{i}$. In stage 2, only predicate entity prediction loss is applied on $R_{i}$, which can be further decomposed into L1 and GIOU losses for object/subject/predicate localization [46] and cross-entropy loss for object/subject/predicate classification. In learning scenario S1, we skip Stage 1, and directly load pre-trained weights of DETR for object detection on the entire training set of Visual Genome [28]. In stage 2 of S1, we freeze the feature extractor, and fine-tune the rest parts of SGTR for predicate entity predictions. As only relationship classes are incrementally introduced in S1, we freeze the entire weights of DETR for detecting all the objects in the scene over tasks. However, empirical results suggest that fine-tuning transformer-based encoders in DETR helps downstream predicate predictions [28]. Even with fine-tuning DETR in S1, we verify that there is minimal forgetting of detecting all objects in the scene over tasks (see Sec. S4 and Sec. 0.A.4.2). Thus, the forgetting observed in S1 could only be attributed to incremental relationship learning. In Stage 1 of S2 and S3 where object classes are also incrementally introduced over tasks, we load weights of the feature extractor, pre-trained on ImageNet [12], and fine-tune the entire DETR [6] over all the tasks. Stage 2 of S2 and S3 is the same as S1.
@@ 0.A.2.1 SGTR Backbone
0.A.2.2 TCNN Backbone
As for TCNN, it employs Faster-RCNN [17] to generate object proposals from a scene image ($I_{k}$). The model extracts visual features for nodes and edges from these proposals. Through message passing, both edge and node GRUs output a structured scene graph. Experimental results are based on the average over three runs.
Given a scene image $I_{i}$, TCNN utilizes Faster-RCNN[17] to generate a set of object proposals. The model subsequently extracts visual features of nodes and edges from the set of object proposals. Finally, both edge and node GRUs output a structured scene graph via message passing. The TCNN is trained in two stages in a supervised manner. In stage 1, only object detection losses in Faster-RCNN[45] are applied on $O_{i}$. We use the cross entropy loss for the object class and $L1$ loss for the bounding box offsets. In stage 2, the visual feature extractor (VGG-16[53] pre-trained on ImageNet [12]) and GRUs layers are trained to predict the final object classes, bounding boxes, and relationship predicates using cross-entropy loss and $L1$ loss. In Learning Scenario 1 (S1), similar to the implementation details of SGTR in Sec. 0.A.2.1, we skip Stage 1, and directly load pre-trained weights of Faster-RCNN for object detection on the entire training set of Visual Genome [28]. In stage 2 of S1, we load the pre-trained weights of the visual feature extractor (pre-trained on ImageNet) and fine-tune the rest parts of the model. In stage 2 of S1, we load the pre-trained weights of visual feature extractor (pre-trained on ImageNet) and fine-tune the rest parts of the model. In Stage 1 of S2 and S3 where object classes are also incrementally introduced over tasks, we load weights of the Faster-RCNN, pre-trained on ImageNet [12], and fine-tune it over all the tasks. Stage 2 of S2 and S3 follows the same training regimes as Stage 2 of S1.
@@ 0.A.2.2 TCNN Backbone
0.A.3 Evaluation Metrics
To assess the catastrophic forgetting of CSEGG models, we define Forgetfullness (F@K) as the difference in R@K on $D_{t=1}$ between the CSEGG models trained at task $t$ and task 1. An ideal CSEGG model could maintain the same $R@K$ on $D_{t=1}$ over tasks; thus, $F=0$ for all tasks. The more negative F is, the more severe in forgetting an model gets.
To assess the overall recall of CSEGG models over tasks, we also report the continual average recall (Avg. R@K). Avg. R@K is computed as the average recall on all the data at the previous and current tasks $D_{i}$, where $i\in\{1,2,...,t\}$.